OpenStudy (anonymous):

Solve the equation by using substitution. x+5y=3 3x-2y=-8

OpenStudy (noseboy908):

Do you know how you would start this, or are you lost already?

OpenStudy (anonymous):

lost already

OpenStudy (noseboy908):

Alright, no problem. Firstly, you'll need to solve for a variable. For simplicity, I'd choose the x in the top equation. Or, if you'd like, pick another. Either way, pick one. We're going to solve for it.

OpenStudy (anonymous):

Ok x is fine so then we would have to get x on one side so would it be x=-5y+3 ?

OpenStudy (noseboy908):

Yes! So, now you have "x=-5y+3", and "3x-2y=-8". Now you'll need to literally substitute the given formula that equals x in the first equation in for the x of the second. If you're following what I mean, rewrite the new second equation as it would appear exactly one step ahead of now.

OpenStudy (anonymous):

3(-5y+3)-2y=-8

OpenStudy (noseboy908):

Correct! Now, using PEMDAS, what's your next step? (Tell me if you reach a point where you understand what you have to do from that point, and I'll check your answer)

OpenStudy (anonymous):

-15y+9-2Y=-8 -17y+9=-8 17y=-17 y=-1

OpenStudy (noseboy908):

Almost completely correct, but you made one minor error: in the third step, the negative on the "17y" disappeared. It should still be there.

OpenStudy (anonymous):

oops so y=1

OpenStudy (noseboy908):

yep. There ya go!

OpenStudy (anonymous):

Perfect thank you!

OpenStudy (noseboy908):

You're most welcome!